# Leveraging Prototypical Representations for Mitigating Social Bias without Demographic Information

### v2

@@ 3.1 Tasks
We use the DIAL Dataset Blodgett et al. (2016), which contains 100K Twitter messages labeled with positive or negative sentiment. The task entails classifying the sentiment expressed in tweets. As a proxy for the writer’s racial identity, each tweet is associated with the sociolect: African American English (AAE) or Standard American English (SAE).
@@ 3.3 Metrics
Method Accuracy $\uparrow$ TPR-GAP $\downarrow$ Accuracy $\uparrow$ TPR-GAP $\downarrow$
JTT $81.34\pm{\scriptstyle 0.81}$ $14.19\pm{\scriptstyle 1.08}$ $78.08\pm{\scriptstyle 1.21}$ $23.67\pm{\scriptstyle 1.87}$
BLIND $82.52\pm{\scriptstyle 0.23}$ $13.76\pm{\scriptstyle 1.18}$ $76.45\pm{\scriptstyle 0.67}$ $22.68\pm{\scriptstyle 2.40}$
@@ Appendix B Fairness Metrics
This metric measures the statistical dependence between the model’s prediction and protected attributes. It employs the Kullback–Leibler divergence between two distributions, namely $KL(P(y),P(\hat{y}|z=z))$, for $z$in $Z$. The sum over $z$ yields a single value describing the model’s independence, irrespective of gold labels. It assesses how the model’s behavior varies across different demographics.
@@
This metric measures the statistical dependence between the target label given the model’s prediction and the protected attributes. It employs $KL(P(y|\hat{y}=\hat{y}),P(y|\hat{y}=\hat{y},z=z))$ for $\hat{y}$in $Y$, $z$in $Z$. The sum over $\hat{y}$ and $z$ results in a single value. Linked to calibration in classification and precision gap, it intuitively assesses whether a model disproportionately promotes or penalizes specific demographic groups.
@@ Appendix C Full Results
Method Accuracy $\uparrow$ TPR-GAP $\downarrow$ Indep $\downarrow$ Sep $\downarrow$ Suff $\downarrow$
Original $83.43\pm{\scriptstyle 0.08}$ $14.66\pm{\scriptstyle 0.51}$ $0.16\pm{\scriptstyle 0.002}$ $2.59\pm{\scriptstyle 0.07}$ $2.46\pm{\scriptstyle 0.08}$
JTT $81.34\pm{\scriptstyle 0.81}$ $14.19\pm{\scriptstyle 1.08}$ $0.16\pm{\scriptstyle 0.002}$ $2.77\pm{\scriptstyle 0.01}$ $2.22\pm{\scriptstyle 0.14}$
BLIND $82.52\pm{\scriptstyle 0.23}$ $13.76\pm{\scriptstyle 1.18}$ $0.16\pm{\scriptstyle 0.002}$ $2.35\pm{\scriptstyle 0.10}$ $2.21\pm{\scriptstyle 0.09}$
@@
Method Accuracy $\uparrow$ TPR-GAP $\downarrow$ Indep $\downarrow$ Sep $\downarrow$ Suff $\downarrow$
Original $79.18\pm{\scriptstyle 0.22}$ $25.34\pm{\scriptstyle 0.82}$ $0.17\pm{\scriptstyle 0.003}$ $0.11\pm{\scriptstyle 0.003}$ $0.08\pm{\scriptstyle 0.004}$
JTT $78.08\pm{\scriptstyle 1.21}$ $23.67\pm{\scriptstyle 1.87}$ $0.18\pm{\scriptstyle 0.01}$ $0.11\pm{\scriptstyle 0.008}$ $0.08\pm{\scriptstyle 0.006}$
BLIND $76.45\pm{\scriptstyle 0.67}$ $22.68\pm{\scriptstyle 2.40}$ $0.14\pm{\scriptstyle 0.001}$ $0.05\pm{\scriptstyle 0.001}$ $0.06\pm{\scriptstyle 0.001}$
@@
Method Accuracy $\uparrow$ TPR-GAP $\downarrow$ Indep $\downarrow$ Sep $\downarrow$ Suff $\downarrow$
Original $83.42\pm{\scriptstyle 0.26}$ $14.60\pm{\scriptstyle 0.81}$ $0.17\pm{\scriptstyle 0.003}$ $2.67\pm{\scriptstyle 0.10}$ $2.51\pm{\scriptstyle 0.09}$
JTT $81.98\pm{\scriptstyle 1.70}$ $14.30\pm{\scriptstyle 0.55}$ $0.18\pm{\scriptstyle 0.004}$ $2.65\pm{\scriptstyle 0.04}$ $2.38\pm{\scriptstyle 0.12}$
BLIND $82.41\pm{\scriptstyle 0.55}$ $13.86\pm{\scriptstyle 1.49}$ $0.16\pm{\scriptstyle 0.003}$ $2.31\pm{\scriptstyle 0.12}$ $2.29\pm{\scriptstyle 0.10}$
@@
Method Accuracy $\uparrow$ TPR-GAP $\downarrow$ Indep $\downarrow$ Sep $\downarrow$ Suff $\downarrow$
Original $78.36\pm{\scriptstyle 0.93}$ $29.62\pm{\scriptstyle 1.58}$ $0.19\pm{\scriptstyle 0.003}$ $0.12\pm{\scriptstyle 0.004}$ $0.09\pm{\scriptstyle 0.005}$
JTT $77.41\pm{\scriptstyle 1.76}$ $28.06\pm{\scriptstyle 2.03}$ $0.18\pm{\scriptstyle 0.013}$ $0.13\pm{\scriptstyle 0.009}$ $0.08\pm{\scriptstyle 0.005}$
BLIND $77.20\pm{\scriptstyle 1.22}$ $27.32\pm{\scriptstyle 2.84}$ $0.15\pm{\scriptstyle 0.001}$ $0.07\pm{\scriptstyle 0.001}$ $0.08\pm{\scriptstyle 0.001}$
@@ C.1 Accuracy Figures
This is a biography about a male. This is a biography about a female.
A man who excelled in his field. A woman who excelled in her field.
He is known for his achievements in various industries. She is known for her achievements in various industries.
A prominent male figure in history. A prominent female figure in history.
His career and accomplishments are well-regarded. Her career and accomplishments are well-regarded.
This biography focuses on the life of a distinguished man. This biography focuses on the life of a distinguished woman.
An influential male individual. An influential female individual.
He made significant contributions to his profession. She made significant contributions to her profession.
This is a story about a man who shaped his industry. This is a story about a woman who shaped her industry.
His impact on his field is noteworthy. Her impact on her field is noteworthy.
@@
This tweet reflects a [sentiment] from a white writer. This tweet reflects a [sentiment] sentiment from a black writer.
A tweet expressing a [sentiment] moment by a white individual. A tweet expressing a [sentiment] moment by a black individual.
A [sentiment] viewpoint shared by a writer using Standart American English. A [sentiment] viewpoint shared by a writer using African American English.
This post, written in standard English, conveys [sentiment] from a white perspective. This post, written in AAE, conveys [sentiment] from a black perspective.
A message filled with [sentiment] from a white communicator. A message filled with [sentiment] from a black communicator.
A white person shares their [sentiment] thoughts in this tweet. A black person shares their [sentiment] thoughts in this tweet.
This is an example of a tweet with [sentiment] in white sociolect. This is an example of a tweet with [sentiment] sentiment in AAE.
A tweet written by a white speaker that conveys [sentiment]. A tweet written by a black speaker that conveys [sentiment].
This post by a white individual radiates [sentiment] and [sentiment]. This post by a black individual radiates [sentiment] and [sentiment].
A [sentiment] perspective presented by a writer using white sociolect. A [sentiment] perspective presented by a writer using African American English.

### v3

@@ 3.1 Tasks
We follow the setup of Elazar and Goldberg (2018), who leveraged a Twitter dataset originally gathered by Blodgett et al. (2016). Elazar and Goldberg (2018) used emojis in the tweets to derive sentiment labels for the classification task. Tweets are labeled with sociolects—African American English (AAE) or Standard American English (SAE)—based on the author’s geo-location, serving as a proxy for their racial identity. We work with a subset of 100K samples, consistent with Orgad and Belinkov (2023).
@@ 3.3 Metrics
Method
Accuracy $\uparrow$ TPR-GAP $\downarrow$ Accuracy $\uparrow$ TPR-GAP $\downarrow$
Original
$83.43\pm{\scriptstyle 0.08}$ $14.66\pm{\scriptstyle 0.51}$ $79.18\pm{\scriptstyle 0.22}$ $25.34\pm{\scriptstyle 0.82}$
JTT
$81.34\pm{\scriptstyle 0.81}$ $14.19\pm{\scriptstyle 1.08}$ $78.08\pm{\scriptstyle 1.21}$ $23.67\pm{\scriptstyle 1.87}$
BLIND
$82.52\pm{\scriptstyle 0.23}$ $13.76\pm{\scriptstyle 1.18}$ $76.45\pm{\scriptstyle 0.67}$ $22.68\pm{\scriptstyle 2.40}$
@@ Appendix B Fairness Metrics
This metric measures the statistical dependence between the model’s prediction and protected attributes. It employs the Kullback–Leibler divergence between two distributions, namely $KL(P(y),P(\hat{y}|z=z))$, for $z$in $Z$. The sum over $z$ yields a single value describing the model’s independence. It assesses how the model’s behavior varies across different demographics.
@@
This metric measures the statistical dependence between the target label given the model’s prediction and the protected attributes. It employs $KL(P(y|\hat{y}=\hat{y}),P(y|\hat{y}=\hat{y},z=z))$ for $\hat{y}$in $Y$, $z$in $Z$. The sum over $\hat{y}$ and $z$ results in a single value. It intuitively assesses whether a model disproportionately promotes or penalizes specific demographic groups.
@@ Appendix C Full Results
Method
Accuracy $\uparrow$ TPR-GAP $\downarrow$ Indep $\downarrow$ Sep $\downarrow$ Suff $\downarrow$
Original
$83.43\pm{\scriptstyle 0.08}$ $14.66\pm{\scriptstyle 0.51}$ $0.16\pm{\scriptstyle 0.002}$ $2.59\pm{\scriptstyle 0.07}$ $2.46\pm{\scriptstyle 0.08}$
JTT
$81.34\pm{\scriptstyle 0.81}$ $14.19\pm{\scriptstyle 1.08}$ $0.16\pm{\scriptstyle 0.002}$ $2.77\pm{\scriptstyle 0.01}$ $2.22\pm{\scriptstyle 0.14}$
BLIND
$82.52\pm{\scriptstyle 0.23}$ $13.76\pm{\scriptstyle 1.18}$ $0.16\pm{\scriptstyle 0.002}$ $2.35\pm{\scriptstyle 0.10}$ $2.21\pm{\scriptstyle 0.09}$
@@
Method
Accuracy $\uparrow$ TPR-GAP $\downarrow$ Indep $\downarrow$ Sep $\downarrow$ Suff $\downarrow$
Original
$79.18\pm{\scriptstyle 0.22}$ $25.34\pm{\scriptstyle 0.82}$ $0.17\pm{\scriptstyle 0.003}$ $0.11\pm{\scriptstyle 0.003}$ $0.08\pm{\scriptstyle 0.004}$
JTT
$78.08\pm{\scriptstyle 1.21}$ $23.67\pm{\scriptstyle 1.87}$ $0.18\pm{\scriptstyle 0.01}$ $0.11\pm{\scriptstyle 0.008}$ $0.08\pm{\scriptstyle 0.006}$
BLIND
$76.45\pm{\scriptstyle 0.67}$ $22.68\pm{\scriptstyle 2.40}$ $0.14\pm{\scriptstyle 0.001}$ $0.05\pm{\scriptstyle 0.001}$ $0.06\pm{\scriptstyle 0.001}$
@@
Method
Accuracy $\uparrow$ TPR-GAP $\downarrow$ Indep $\downarrow$ Sep $\downarrow$ Suff $\downarrow$
Original
$83.42\pm{\scriptstyle 0.26}$ $14.60\pm{\scriptstyle 0.81}$ $0.17\pm{\scriptstyle 0.003}$ $2.67\pm{\scriptstyle 0.10}$ $2.51\pm{\scriptstyle 0.09}$
JTT
$81.98\pm{\scriptstyle 1.70}$ $14.30\pm{\scriptstyle 0.55}$ $0.18\pm{\scriptstyle 0.004}$ $2.65\pm{\scriptstyle 0.04}$ $2.38\pm{\scriptstyle 0.12}$
BLIND
$82.41\pm{\scriptstyle 0.55}$ $13.86\pm{\scriptstyle 1.49}$ $0.16\pm{\scriptstyle 0.003}$ $2.31\pm{\scriptstyle 0.12}$ $2.29\pm{\scriptstyle 0.10}$
@@
Method
Accuracy $\uparrow$ TPR-GAP $\downarrow$ Indep $\downarrow$ Sep $\downarrow$ Suff $\downarrow$
Original
$78.36\pm{\scriptstyle 0.93}$ $29.62\pm{\scriptstyle 1.58}$ $0.19\pm{\scriptstyle 0.003}$ $0.12\pm{\scriptstyle 0.004}$ $0.09\pm{\scriptstyle 0.005}$
JTT
$77.41\pm{\scriptstyle 1.76}$ $28.06\pm{\scriptstyle 2.03}$ $0.18\pm{\scriptstyle 0.013}$ $0.13\pm{\scriptstyle 0.009}$ $0.08\pm{\scriptstyle 0.005}$
BLIND
$77.20\pm{\scriptstyle 1.22}$ $27.32\pm{\scriptstyle 2.84}$ $0.15\pm{\scriptstyle 0.001}$ $0.07\pm{\scriptstyle 0.001}$ $0.08\pm{\scriptstyle 0.001}$
@@ C.1 Accuracy Figures
In Figure 3 we provide the accuracy of different methods across varying dataset sizes.
@@
This is a biography about a male.
This is a biography about a female.
A man who excelled in his field.
A woman who excelled in her field.
He is known for his achievements in various industries.
She is known for her achievements in various industries.
A prominent male figure in history.
A prominent female figure in history.
His career and accomplishments are well-regarded.
Her career and accomplishments are well-regarded.
This biography focuses on the life of a distinguished man.
This biography focuses on the life of a distinguished woman.
An influential male individual.
An influential female individual.
He made significant contributions to his profession.
She made significant contributions to her profession.
This is a story about a man who shaped his industry.
This is a story about a woman who shaped her industry.
His impact on his field is noteworthy.
Her impact on her field is noteworthy.
@@
This tweet reflects a [sentiment] from a white writer.
This tweet reflects a [sentiment] sentiment from a black writer.
A tweet expressing a [sentiment] moment by a white individual.
A tweet expressing a [sentiment] moment by a black individual.
A [sentiment] viewpoint shared by a writer using Standart American English.
A [sentiment] viewpoint shared by a writer using African American English.
This post, written in standard English, conveys [sentiment] from a white perspective.
This post, written in AAE, conveys [sentiment] from a black perspective.
A message filled with [sentiment] from a white communicator.
A message filled with [sentiment] from a black communicator.
A white person shares their [sentiment] thoughts in this tweet.
A black person shares their [sentiment] thoughts in this tweet.
This is an example of a tweet with [sentiment] in white sociolect.
This is an example of a tweet with [sentiment] sentiment in AAE.
A tweet written by a white speaker that conveys [sentiment].
A tweet written by a black speaker that conveys [sentiment].
This post by a white individual radiates [sentiment] and [sentiment].
This post by a black individual radiates [sentiment] and [sentiment].
A [sentiment] perspective presented by a writer using white sociolect.
A [sentiment] perspective presented by a writer using African American English.